# Multi-stage tomography based on eigenanalysis for high-dimensional dense unitary quantum processes

### v1

@@ III.4 Third part of the methods
Besides, $d_{2}$, i.e. the number of different eigenvalues of the input density matrix (as defined in Section III.2), is upper bounded by the value $n_{max}$, as explained above. Therefore, due to (26) and (19), the state space dimension $d$ that can be handled by the methods proposed in the present section is upper bounded by:
@@
When $d_{2}$ is moreover set to its maximum value $n_{max}$, we obtain $d=(n_{max})^{2}$, i.e. we reach the upper bound defined by (27).
@@
III.5 Fourth part of the methods
The properties of $\widehat{U}_{4}$ defined at the end of the previous section are exactly the same as those obtained in our single-stage methods for $\widehat{U}_{2}$: see Section II.1, including (6). Therefore, the final stage of the method considered here consists of processing $\widehat{U}_{4}$ in the same way as we processed $\widehat{U}_{2}$ in Section II.2. This yields a matrix $\widehat{U}_{5}$ that here again succeeds in restoring $U$, up to only a global phase factor.
A pseudo-code of the version of our algorithm described in this Section III is provided in Algorithm 2. This algorithm is called EQPT2. Two variants (called EQPT3 and EQPT4) of two-stage QPT methods are moreover presented in Appendix E.
Input : a) Estimate $\widehat{\rho}_{2}$ of output density matrix $\rho_{2}$ (provided by QST and obtained for input density matrix (20) with the diagonal values defined in (21)). b) Estimate $\widehat{\rho}_{6}$ of output density matrix $\rho_{6}$ (provided by QST and obtained for input density matrix (23) with the diagonal values defined in (21)). c) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
@@ III.5 Fourth part of the methods
       /* [see Sect. III.2] If this is not performed by the considered QST algorithm, then derive Hermitian and unit-trace version $\widehat{\rho}_{4}$ of $\widehat{\rho}_{2}$ as follows (otherwise, $\widehat{\rho}_{4}=\widehat{\rho}_{2}$): */
       Compute $\widehat{\rho}_{3}$ and $\widehat{\rho}_{4}$ by using (2) and (3);
@@
       /* [see Sect. III.3] If this is not performed by the considered QST algorithm, then derive Hermitian and unit-trace version $\widehat{\rho}_{8}$ of $\widehat{\rho}_{6}$ as follows (otherwise, $\widehat{\rho}_{8}=\widehat{\rho}_{6}$): */
       Compute $\widehat{\rho}_{7}$ and $\widehat{\rho}_{8}$ as in (2) and (3);
@@
       /* [see Sect. III.4] Combine $\widehat{U}_{2}$ and $\widehat{U}_{3}$/
@@
                   Set column of $\widehat{U}_{4}$ whose index is defined by (30): set it to first vector provided by CCA of 1) the set of columns of $\widehat{U}_{2}$ with indices defined by (22) and 2) the set of columns of $\widehat{U}_{3}$ with indices defined by (24);
@@
      /* [see Sect. III.5] If the above QST algorithm does not yield a unit-norm ket $\lvert\widehat{\Psi}_{2}\rangle$, then create it as follows: */
       Divide $\lvert\widehat{\Psi}_{2}\rangle$ by its norm;
@@
       Compute $\lvert\Psi_{3}\rangle$ by using (12) with $\widehat{U}_{2}$ replaced by $\widehat{U}_{4}$;
       Compute $\widehat{U}_{5}$ by using (16) with $\widehat{U}_{2}$ replaced by $\widehat{U}_{4}$;
@@
Algorithm 2 EQPT2: second Eigenanalysis-based QPT algorithm (composed of two stages).
@@ IV Proposed multi-stage QPT methods
Moreover, these blocks are organized as follows: from the top to the bottom of the diagonal of $\rho_{1}$, we have a block of values equal to $r_{1}$, then a block of values $r_{2}$, and so on until a block of values $r_{d_{2}}$, and then the same structure again and again, i.e. blocks of values $r_{1},r_{2},\dots,r_{d_{2}}$ (when there exist several blocks containing the same value, i.e. when $b_{n}>1$). Here again, we request these possible values to be defined in decreasing order, i.e. $r_{1}>r_{2}>\dots>r_{d_{2}}$, and to take the same values as in (21). The two extreme cases of this general framework are nothing but the two specific cases that we already used for the above-defined two-stage methods: when setting ${b_{\ell}}=0$ and hence $b_{n}=1$ and $b_{s}=d_{1}$, we obtain the density matrix of (20) whereas with ${b_{\ell}}=\log_{2}d_{1}$ and hence $b_{n}=d_{1}$ and $b_{s}=1$, we get the density matrix of (23).
@@
The main idea of the above-defined two-stage methods was to exploit intersections of two subspaces of eigendecompositions resulting from the above density matrices (20) and (23). Here, we extend this idea to general multi-stage methods, with $(\log_{2}d_{1}+1)$ stages. To this end, we consider all $(\log_{2}d_{1}+1)$ input density matrices that belong to the above-defined general framework, by considering all $(\log_{2}d_{1}+1)$ possible values of ${b_{\ell}}$, i.e. all integers from 0 to $\log_{2}d_{1}$. Using the same approach as in Section III (i.e. performing the eigendecomposition and reordering for the estimated output density matrix), for each value of ${b_{\ell}}$ and hence of $b_{n}$ and $b_{s}$, we obtain a set of $d_{2}$ subspaces denoted as ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$, with $m_{{b_{\ell}}}\in\{1,\dots,d_{2}\}$ (each index $m_{{b_{\ell}}}$ is used to define each subspace in the eigendecomposition of the stage ${b_{\ell}}$ of the algorithm proposed here; its correspondence with the structure of the input density matrix is not the same as when we used the indices $m_{1}$ and $m_{2}$ in Section III). These subspaces ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$ are defined in more detail in Appendix F. We then extend the approach of Section III by performing subspace intersections as follows. We successively consider all $d_{2}^{(\log_{2}d_{1}+1)}$ possible values of the set of $(\log_{2}d_{1}+1)$ subspace indices $m_{{b_{\ell}}}$ (with $m_{{b_{\ell}}}\in\{1,\dots,d_{2}\}$), i.e. we consider all possible values of the set of indices $\{m_{0},\dots,m_{\log_{2}d_{1}}\}$. For each suh set of indices, we derive the intersection of the corresponding $(\log_{2}d_{1}+1)$ subspaces ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$. We thus form $d_{2}^{(\log_{2}d_{1}+1)}$ subspace intersections. As in Section III.4, we want each of these intersections to be one-dimensional (or possibly to be restricted to the null vector for some of these intersections). In order to identify all $d$ columns of $U$, the number of subspace intersections should be at least equal to $d$, i.e. using (19)
$d_{2}^{\log_{2}d_{1}}d_{2}\geq d_{1}d_{2}$ (33)
i.e., taking the logarithm, in base 2, of that equation:
@@
In Appendix F, we prove that, for $d_{2}=2$, the above-defined subspace intersections are each indeed restricted to one dimension and yield all columns of $U$ (up to scale or phase factors). Moreover, we here illustrate this property and thus the proposed complete multi-stage methods by considering the case $d=8$, $d_{2}=2$ and hence $d_{1}=4$ and $(\log_{2}d_{1}+1)=3$ stages. During the first stage, that corresponds to ${b_{\ell}}=0$, the eigendecomposition allows one to identify two subspaces ${\cal S}_{0,m_{0}}$, respectively with $m_{0}=1$ and $m_{0}=2$, and that respectively correspond to the following sets of 4 indices of columns of $U$: $\{1,2,3,4\}$ and $\{5,6,7,8\}$. Then, during the second stage, that corresponds to ${b_{\ell}}=1$, the eigendecomposition allows one to identify two subspaces ${\cal S}_{1,m_{1}}$, respectively with $m_{1}=1$ and $m_{1}=2$, and that respectively correspond to the following sets of 4 indices of columns of $U$: $\{1,2,5,6\}$ and $\{3,4,7,8\}$. Finally, the third stage, that corresponds to ${b_{\ell}}=2$, yields the following sets of 4 indices of columns of $U$: $\{1,3,5,7\}$ and $\{2,4,6,8\}$. We then derive the corresponding subspace intersections, first considering ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}$ as intermediate quantities, and finally deriving ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}\cap{\cal S}_{2,m_{2}}$. All these intersections are listed in Table 1. This proves that the proposed approach yields 8 intersections that are each restricted to one column of $U$ and that this set of intersections yields all the individual columns of $U$, as required.
$\{1,2,3,4\}$$\{1,2,5,6\}$$\{1,3,5,7\}$$\{1,2\}$$\{1\}$
$\{1,2,3,4\}$$\{1,2,5,6\}$$\{2,4,6,8\}$$\{1,2\}$$\{2\}$
$\{1,2,3,4\}$$\{3,4,7,8\}$$\{1,3,5,7\}$$\{3,4\}$$\{3\}$
$\{1,2,3,4\}$$\{3,4,7,8\}$$\{2,4,6,8\}$$\{3,4\}$$\{4\}$
$\{5,6,7,8\}$$\{1,2,5,6\}$$\{1,3,5,7\}$$\{5,6\}$$\{5\}$
$\{5,6,7,8\}$$\{1,2,5,6\}$$\{2,4,6,8\}$$\{5,6\}$$\{6\}$
$\{5,6,7,8\}$$\{3,4,7,8\}$$\{1,3,5,7\}$$\{7,8\}$$\{7\}$
$\{5,6,7,8\}$$\{3,4,7,8\}$$\{2,4,6,8\}$$\{7,8\}$$\{8\}$
Table 1: Multi-stage QPT method. Each column of the table contains the indices of the column vectors that correspond to the considered subspace or subspace intersection defined in the first row of the table.
Appendix F also outlines a possible structure for the pseudo-code of that multi-stage QPT method for an arbitrary state space dimension.
@@ V Test results
More precisely, to model a ket estimate (e.g. for the ket $\lvert\Psi_{2}\rangle$ of Section II.2), we compute its theoretical value in the considered conditions (e.g. using (9) with the input ket defined before (15) and a known process matrix $U$) and we add random, independent, complex-valued fluctuations to all components of this ket. Their real and imaginary parts are separately created with a random generator that has a probability density function that is uniform over the range $[-w/2,w/2]$, where the parameter $w$ defines the magnitude of the modeled estimation fluctuations (their variance is thus equal to $w^{2}/12$). This parameter is hereafter varied so as to investigate the sensitivity of the proposed QPT methods to the magnitude of QST estimation errors. This simple error model has the advantage of being easily interpretable. The error model for density matrix estimation is defined accordingly and is presented in Appendix G.
Each elementary test uses a randomly drawn $d\times d$ process matrix $U$, obtained as follows. We first create a $d\times d$ real-valued matrix with each element randomly, uniformly, drawn in the interval $[0,1]$. We then apply the so-called “QR algorithm” (this standard algorithm is e.g. defined in [43] p. 112) to the above-defined matrix. This especially yields the so-called “Q factor” of that QR decomposition. This factor is guaranteed to be a unitary matrix (see [43] p. 112) and we therefore use it as our matrix $U$.
@@
Figure 1: Normalized Root Mean Square Error (NRMSE) of estimation of process matrix $U$ of proposed Quantum Process Tomography (QPT) methods EQPT1, EQPT2 and EQPT3, vs. dimension $d$ of state space, for several values of the parameter $w$ that defines the magnitude of the estimation errors of Quantum State Tomography (QST).
@@
The reported tests were carried out with a standard PC (Intel(R) Core(TM) i5-7500 CPU running at 3.40 GHz with 16 GB of RAM), using the Matlab software environment. As a first step, Fig. 1 shows the results thus obtained with our single-stage and two-stage methods, namely EQPT1 (black solid lines), EQPT2 (blue dash doted lines) and EQPT3 (red dashed lines). The results for EQPT1 are provided for all values of the state space dimension $d$ that take the form $d=2^{q}$, where $q$ is the equivalent number of qubits, that is here varied from 2 to 13. $q=13$ is the maximum number that was accepted in the considered hardware and software environment without running out of memory (for non-optimized code). The same approach applies to EQPT2 and EQPT3 except that, due to (29), only even values of $q$ were considered. Moreover, for EQPT2 and EQPT3, performance turned out to often be somewhat degraded for the lowest possible state space dimension, namely $d=4$, that corresponds to $q=2$ qubits. This is not an issue, since EQPT2 and EQPT3 were developed to address higher values of $d$ (EQPT1 can be used instead for lower values of $d$). Therefore, for better readability, the results of EQPT2 and EQPT3 for $d=4$ are skipped in Fig. 1. It should also be noted that the complexity of the configurations that could be tested was limited by memory requirements, as stated above, i.e. not by the computational load of the proposed algorithms, that remained reasonable up to $q=13$. For instance, for each of the 100 process matrix estimations, the average execution time used for a) creating data (thus only “modeling” QST instead of running a QST algorithm: see above), b) running the QPT algorithm and c) displaying results, was as follows, respectively for EQPT1, EQPT2 and EQPT3: 1) 0.9 ms, 2.6 ms and 2.8 ms for $q=4$ qubits, 2) 22 ms, 170 ms and 200 ms for $q=8$ qubits and 3) 23 s, 180 s and 230 s for $q=12$ qubits.
@@
The following conlusions can be drawn from Fig. 1, when comparing the proposed methods for any given QST quality, defined by the considered value of $w$. The two-stage methods EQPT2 and EQPT3 yield almost the same performance, with a slight advantage for EQPT3 (as expected from Appendix E), that is more significant for lower values of the space dimension $d$. Moreover, these two-stage methods always yield better and often much better performance than the single-stage method EQPT1, as expected from the beginning of Section III. Up to a certain space dimension $d$, the gap between these two types of methods increases with the space dimension (and associated number $q$ of qubits): whereas all methods yield linear variations of the logarithm of NRMSE with respect to the logarithm of $d$ (i.e. with respect to $q$ itself), the slope of these variations is significantly lower for EQPT2 and EQPT3, and these methods already start with a lower NRMSE than EQPT1 for the lowest values of $d$. Depending on the considered conditions, the EQPT2 and EQPT3 methods decrease the NRMSE by a factor up to 6, and hence the NMSE by a factor up to 36, as compared with EQPT1. Moreover, for the lowest QST error magnitude considered in these tests, for all analyzed state space dimensions, EQPT2 and EQPT3 yield an NRMSE limited to about $10^{-1}$ and hence an NMSE limited to about $10^{-2}$.
@@
Figure 2: Normalized Root Mean Square Error (NRMSE) of estimation of process matrix $U$ of proposed Quantum Process Tomography (QPT) methods EQPT1, EQPT2 and EQPT5, vs. dimension $d$ of state space, for several values of the parameter $w$ that defines the magnitude of the estimation errors of Quantum State Tomography (QST).
@@
We now move to the multi-stage method EQPT5, whose performance is shown in Fig. 2 (red dashed lines), together with the results already provided above for EQPT1 (black solid lines) and EQPT2 (blue dash doted lines). The results for EQPT5 are reported for values of the state space dimension $d$ that take the form $d=2^{q}$, where $q$ is the equivalent number of qubits, varied as follows. For a large $q$, EQPT5 yields a significantly higher computational complexity than the single-stage and two-stage methods, due to the higher number of eigendecompositions it performs and to the associated program structure (recursive function with a binary tree structure, as explained in Appendix F). Unlike for the previous proposed methods, the applicability of EQPT5 to a large number of qubits was therefore more limited by its computational load than by its memory requirements, when using the considered basic hardware and software means. Therefore, to keep a reasonable simulation duration (for all estimations of the process matrix), EQPT5 was only tested up to $q=10$ qubits when performing 100 above-defined elementary tests, and also for $q=11$ qubits but then with only 10 elementary tests. Moreover, as with EQPT2, we hereafter skip the results for a low dimension, that is for $d\leq 4$, which is not the case of interest for a multi-stage method. In the considered range of space dimensions, the above-defined average execution time per process matrix estimation was e.g. as follows for EQPT5: 54 ms for $q=4$ qubits, 44 s for $q=8$ qubits, 490 s $\simeq$ 8 mn for $q=10$ qubits and 7600 s $\simeq$ 2 h 07 mn for $q=11$ qubits.
@@
Fig. 2 yields the following conclusions for EQPT5. For the lowest considered values of the space dimension, EQPT5 starts with a somewhat higher NRMSE than EQPT1 and EQPT2. But, an attractive feature of EQPT5 is that its NRMSE then increases more slowly that those of the other methods, with respect to the space dimension (this again yields linear variations in the considered logarithm representation). EQPT5 thus outperforms the other two methods for higher space dimensions, and the gap between them increases with that dimension, up to a certain dimension. EQPT5 is thus attractive for high space dimensions, so that it actually reaches the goal for which it was designed. Depending on the considered conditions, EQPT5 decreases the NRMSE by a factor up to 3.2, and hence the NMSE by a factor up to more than 10, as compared with EQPT2. Moreover, for the lowest QST error magnitude considered in these tests, for all analyzed state space dimensions (which are slightly lower than those considered above for EQPT1 and EQPT2, however), EQPT5 yield an NRMSE limited to around $3.2\times 10^{-2}$ and hence an NMSE limited to around $10^{-3}$.

### v2

@@ III-D Third part of the methods
Besides, the number $d_{2}$ of different eigenvalues of the input density matrix (as defined in Section III-B) is upper bounded by the value $n_{max}$, as explained above. Therefore, due to (25) and (16), the state space dimension $d$ that can be handled by the methods proposed in the present section is upper bounded by:
@@
When $d_{2}$ is moreover set to its maximum value $n_{max}$, we obtain $d=(n_{max})^{2}$, so that we reach the upper bound defined by (26).
@@
III-E Fourth part of the methods
The properties of $\widehat{U}_{4}$ defined at the end of the previous section are exactly the same as those obtained in our single-stage methods for $\widehat{U}_{2}$: see Section II-A, including (4). Therefore, the final stage of the method considered here consists of processing $\widehat{U}_{4}$ in the same way as we processed $\widehat{U}_{2}$ in Section II-B. This yields a matrix $\widehat{U}_{5}$ that here again succeeds in restoring $U$, up to only a global phase factor.
A pseudo-code of the version of our algorithm described in this Section III is provided in Algorithm 2. This algorithm is called EQPT2. Two variants (called EQPT3 and EQPT4) of two-stage QPT methods are moreover presented in Appendix J.
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by QST and obtained for input density matrix (17) with the diagonal values defined in (19)). b) Estimate $\widehat{\rho}_{8}$ of output density matrix $\rho_{6}$ (provided by QST and obtained for input density matrix (21) with the diagonal values defined in (19)). c) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
@@ III-E Fourth part of the methods
       /* Combine $\widehat{U}_{2}$ and $\widehat{U}_{3}$: */
@@
                   Set column of $\widehat{U}_{4}$ with index equal to $((m_{1}-1)d_{1}+m_{2})$: Set it to first vector provided by CCA of 1) the set of columns of $\widehat{U}_{2}$ with indices $((m_{1}-1)d_{1}+n_{1})$ with $n_{1}\in\{1,\dots,d_{1}\}$ and 2) the set of columns of $\widehat{U}_{3}$ with indices $((m_{2}-1)d_{1}+n_{2})$ with $n_{2}\in\{1,\dots,d_{1}\}$;
@@
       $\lvert\Psi_{3}\rangle=\widehat{U}_{4}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}=\widehat{U}_{4}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
@@
Algorithm 2 EQPT2: second Eigenanalysis-based QPT algorithm (composed of two stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
@@ IV Proposed multi-stage QPT methods
Moreover, these blocks are organized as follows: from the top to the bottom of the diagonal of $\rho_{1}$, we have a block of values equal to $r_{1}$, then a block of values $r_{2}$, and so on until a block of values $r_{d_{2}}$, and then the same structure again and again, that is, blocks of values $r_{1},r_{2},\dots,r_{d_{2}}$ (when there exist several blocks containing the same value, i.e. when $b_{n}>1$). Here again, we request these possible values to be defined in decreasing order, i.e. $r_{1}>r_{2}>\dots>r_{d_{2}}$, and to take the same values as in (19). The two extreme cases of this general framework are nothing but the two specific cases that we already used for the above-defined two-stage methods: when setting ${b_{\ell}}=0$ and hence $b_{n}=1$ and $b_{s}=d_{1}$, we obtain the density matrix of (17) whereas with ${b_{\ell}}=\log_{2}d_{1}$ and hence $b_{n}=d_{1}$ and $b_{s}=1$, we get the density matrix of (21).
@@
The main idea of the above-defined two-stage methods was to exploit intersections of two subspaces of eigendecompositions resulting from the above density matrices (17) and (21). Here, we extend this idea to general multi-stage methods, with $(\log_{2}d_{1}+1)$ stages. To this end, we consider all $(\log_{2}d_{1}+1)$ input density matrices that belong to the above-defined general framework, by considering all $(\log_{2}d_{1}+1)$ possible values of ${b_{\ell}}$, i.e. all integers from 0 to $\log_{2}d_{1}$. Using the same approach as in Section III (that is, performing the eigendecomposition and reordering for the estimated output density matrix), for each value of ${b_{\ell}}$ and hence of $b_{n}$ and $b_{s}$, we obtain a set of $d_{2}$ subspaces denoted as ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$, with $m_{{b_{\ell}}}\in\{1,\dots,d_{2}\}$ (each index $m_{{b_{\ell}}}$ is used to define each subspace in the eigendecomposition of the stage ${b_{\ell}}$ of the algorithm proposed here; its correspondence with the structure of the input density matrix is not the same as when we used the indices $m_{1}$ and $m_{2}$ in Section III). These subspaces ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$ are defined in more detail in Appendix K. We then extend the approach of Section III by performing subspace intersections as follows. We successively consider all $d_{2}^{(\log_{2}d_{1}+1)}$ possible values of the set of $(\log_{2}d_{1}+1)$ subspace indices $m_{{b_{\ell}}}$ (with $m_{{b_{\ell}}}\in\{1,\dots,d_{2}\}$), i.e. we consider all possible values of the set of indices $\{m_{0},\dots,m_{\log_{2}d_{1}}\}$. For each such set of indices, we derive the intersection of the corresponding $(\log_{2}d_{1}+1)$ subspaces ${\cal S}_{{b_{\ell}},m_{{b_{\ell}}}}$. We thus form $d_{2}^{(\log_{2}d_{1}+1)}$ subspace intersections. As in Section III-D, we want each of these intersections to be one-dimensional (or possibly to be restricted to the null vector for some of these intersections). In order to identify all $d$ columns of $U$, the number of subspace intersections should be at least equal to $d$. Using (16), this condition reads
$d_{2}^{\log_{2}d_{1}}d_{2}\geq d_{1}d_{2}.$ (32)
Taking the logarithm, in base 2, of that equation yields
@@
In Appendix K-A we prove that, for $d_{2}=2$, the above-defined subspace intersections are each indeed restricted to one dimension and yield all columns of $U$ (up to scale or phase factors). Moreover, we illustrate this property and thus the proposed complete multi-stage methods in Appendix K-B. Finally, Appendix K-C contains a pseudo-code for that multi-stage QPT method applicable to an arbitrary state space dimension and called EQPT5.
@@ V Test results
More precisely, to model a ket estimate (e.g. for the ket $\lvert\Psi_{2}\rangle$ of Section II-B), we compute its theoretical value in the considered conditions (e.g. using (6) with the input ket defined before (12) and a known process matrix $U$) and we add random, independent, complex-valued fluctuations to all components of this ket. Their real and imaginary parts are separately created with a random generator that has a probability density function that is uniform over the range $[-w/2,w/2]$, where the parameter $w$ defines the magnitude of the modeled estimation fluctuations (their variance is thus equal to $w^{2}/12$). This parameter is hereafter varied so as to investigate the sensitivity of the proposed QPT methods to the magnitude of QST estimation errors. This simple error model has the advantage of being easily interpretable. The error model for density matrix estimation is defined accordingly and is presented in Appendix L.
Each elementary test uses a randomly drawn $d\times d$ process matrix $U$, obtained as follows. We first create a $d\times d$ real-valued matrix with each element randomly, uniformly, drawn in the interval $[0,1]$. We then perform a QR decomposition of this matrix (e.g. using one of the algorithms defined in [48] p. 112, [74] p. 223, [75]). This especially yields the so-called “Q factor” of that QR decomposition (see e.g. the software package [75]). This factor is guaranteed to be a unitary matrix (see [48] p. 112, [74] p. 223, [75]) and we therefore use it as our matrix $U$.
@@
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=]FIG1.png Normalized Root Mean Square Error (NRMSE) of estimation of process matrix $U$ of proposed Quantum Process Tomography (QPT) methods EQPT1 (black solid lines), EQPT2 (blue dash doted lines) and EQPT3 (red dashed lines), vs. number $q$ of qubits, for several values of the parameter $w$ that defines the magnitude of the estimation errors of Quantum State Tomography (QST).
@@
The reported tests were carried out with a standard PC (Intel(R) Core(TM) i5-7500 CPU running at 3.40 GHz with 16 GB of RAM), using the Matlab software environment. As a first step, Fig. V shows the results thus obtained with our single-stage and two-stage methods, namely EQPT1 (black solid lines), EQPT2 (blue dash doted lines) and EQPT3 (red dashed lines). The results for EQPT1 are provided for all values of the state space dimension $d$ that take the form $d=2^{q}$, where $q$ is the equivalent number of qubits, that is here varied from 2 to 13. $q=13$ is the maximum number that was accepted in the considered hardware and software environment without running out of memory (for non-optimized code). The same approach applies to EQPT2 and EQPT3 except that, due to (28), only even values of $q$ were considered. Moreover, for EQPT2 and EQPT3, performance turned out to often be somewhat degraded for the lowest possible state space dimension, namely $d=4$, that corresponds to $q=2$ qubits. This is not an issue, since EQPT2 and EQPT3 were developed to address higher values of $d$ (EQPT1 can be used instead for lower values of $d$). Therefore, for better readability, the results of EQPT2 and EQPT3 for $q=2$ are skipped in Fig. V. It should also be noted that the complexity of the configurations that could be tested was limited by memory requirements, as stated above, hence not by the computational load of the proposed algorithms, that remained reasonable up to $q=13$. For instance, for each of the 100 process matrix estimations, the average execution time used for a) creating data (thus only “modeling” QST instead of running a QST algorithm: see above), b) running the QPT algorithm and c) displaying results, was as follows, respectively for EQPT1, EQPT2 and EQPT3: 1) $\simeq 1$ ms, 2.6 ms and 2.8 ms for $q=4$ qubits, 2) 22 ms, 170 ms and 200 ms for $q=8$ qubits and 3) 23 s, 180 s and 230 s for $q=12$ qubits. The execution times for the other numbers of qubits are provided in Table I.
Method number of qubits
1 2 3 4 5 6 7 8 9 10 11 12 13
EQPT1 $\simeq 1$ ms $\simeq 1$ ms $\simeq 1$ ms $\simeq 1$ ms 2.2 ms 3.9 ms 7.2 ms 22 ms 68 ms 400 ms 3.1 s 23 s 170 s
EQPT2 2.6 ms 15 ms 170 ms 4.3 s 180 s
EQPT3 2.8 ms 18 ms 200 ms 5.0 s 230 s
EQPT5 2.3 ms 5.4 ms 27 ms 100 ms 580 ms 4.4 s 38 s 490 s 7600 s
$\simeq$ 8 mn $\simeq$ 2 h 07 mn
TABLE I: Execution times (averaged over tests with all values of $w$) of the EQPT1, EQPT2, EQPT3 and EQPT5 methods vs. number of qubits.
@@
The following conlusions can be drawn from Fig. V, when comparing the proposed methods for any given QST quality, defined by the considered value of $w$. The two-stage methods EQPT2 and EQPT3 yield almost the same performance, with a slight advantage for EQPT3 (as expected from Appendix J), that is more significant for lower values of the space dimension $d$. Moreover, these two-stage methods always yield better and often much better performance than the single-stage method EQPT1, as expected from the beginning of Section III. Up to a certain space dimension $d$, the gap between these two types of methods increases with the space dimension and hence with the associated number $q$ of qubits: whereas all methods yield linear variations of the logarithm of NRMSE with respect to $q$ (i.e. logarithm of $d$), the slope of these variations is significantly lower for EQPT2 and EQPT3, and these methods already start with a lower NRMSE than EQPT1 for the lowest values of $d$. Depending on the considered conditions, the EQPT2 and EQPT3 methods decrease the NRMSE by a factor up to 6, and hence the NMSE by a factor up to 36, as compared with EQPT1. Moreover, for the lowest QST error magnitude considered in these tests, for all analyzed state space dimensions, EQPT2 and EQPT3 yield an NRMSE limited to about $10^{-1}$ and hence an NMSE limited to about $10^{-2}$.
[…]
@@
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=]FIG2.png Normalized Root Mean Square Error (NRMSE) of estimation of process matrix $U$ of proposed Quantum Process Tomography (QPT) methods EQPT1 (black solid lines), EQPT2 (blue dash doted lines) and EQPT5 (red dashed lines), vs. (number $q$ of qubits) for several values of the parameter $w$ that defines the magnitude of the estimation errors of Quantum State Tomography (QST).
@@
We now move to the multi-stage method EQPT5, whose performance is shown in Fig. I (red dashed lines), together with the results already provided above for EQPT1 (black solid lines) and EQPT2 (blue dash doted lines). The results for EQPT5 are reported for values of the state space dimension $d$ that take the form $d=2^{q}$, where $q$ is the equivalent number of qubits, varied as follows. For a large $q$, EQPT5 yields a significantly higher computational complexity than the single-stage and two-stage methods, due to the higher number of eigendecompositions it performs and to the associated program structure (recursive function with a binary tree structure, as explained in Appendix K). Unlike for the previous proposed methods, the applicability of EQPT5 to a large number of qubits was therefore more limited by its computational load than by its memory requirements, when using the considered basic hardware and software means. Therefore, to keep a reasonable simulation duration (for all estimations of the process matrix), EQPT5 was only tested up to $q=10$ qubits when performing 100 above-defined elementary tests, and also for $q=11$ qubits but then with only 10 elementary tests. Moreover, as with EQPT2, we hereafter skip the results for a low dimension, that is for $d\leq 4$, which is not the case of interest for a multi-stage method. In the considered range of space dimensions, the above-defined average execution time per process matrix estimation was e.g. as follows for EQPT5: 5.4 ms for $q=4$ qubits, 4.4 s for $q=8$ qubits, 490 s $\simeq$ 8 mn for $q=10$ qubits and 7600 s $\simeq$ 2 h 07 mn for $q=11$ qubits. The execution times for the other numbers of qubits are provided in Table I.
@@
Fig. I yields the following conclusions for EQPT5. For the lowest considered values of the space dimension, EQPT5 starts with a somewhat higher NRMSE than EQPT1 and EQPT2. But, an attractive feature of EQPT5 is that its NRMSE then increases more slowly that those of the other methods, with respect to the space dimension. This again yields linear variations of the logarithm of NRMSE with respect to the number of qubits. EQPT5 thus outperforms the other two methods for higher space dimensions, and the gap between them increases with that dimension, up to a certain dimension. EQPT5 is thus attractive for high space dimensions, so that it actually reaches the goal for which it was designed. Depending on the considered conditions, EQPT5 decreases the NRMSE by a factor up to 3.2, and hence the NMSE by a factor up to more than 10, as compared with EQPT2. Moreover, for the lowest QST error magnitude considered in these tests, for all analyzed state space dimensions (which are slightly lower than those considered above for EQPT1 and EQPT2, however), EQPT5 yield an NRMSE limited to around $3.2\times 10^{-2}$ and hence an NMSE limited to around $10^{-3}$.